OpenStudy (anonymous):

How do you simplify x^21/5

OpenStudy (imstuck):

\[\frac{ x ^{21} }{ 5 }\]Is that it?

OpenStudy (anonymous):

try this: http://www.cymath.com/answer.php

OpenStudy (imstuck):

or is it\[x ^{\frac{ 21 }{ 5 }}\]

OpenStudy (anonymous):

No its x^21/5 its an exponent

OpenStudy (anonymous):

just type the problem and it will solve it or simplfy

OpenStudy (anonymous):

Thanks

geerky42 (geerky42):

Either way, I don't think it can be more simplified.

OpenStudy (anonymous):

cool im glad to help

OpenStudy (anonymous):

gtg later

OpenStudy (anonymous):

Can't think of anything besides this: \[x ^{\frac{ 21 }{ 5 }} = \sqrt[5]{x ^{21}}\]

OpenStudy (anonymous):

\[\huge \sqrt[5]{x^{21}}=x^4\sqrt[5]{x}\]

OpenStudy (anonymous):

I went on that website above but it cant help so I think you cant simplify it